If the cost of Production given by C=50+Q+Q and the demand function is P=40-Q, then what is the profit maximizing level of production for the monopolist

1
Expert's answer
2022-03-27T11:02:35-0400

The marginal cost of production:

"C^{\prime}=(50+2Q)^{\prime}=2"

The profit-maximizing level of production for the monopolist can be find from the equation:

"P*Q=40Q-Q^2=2"

"Q^2-40Q+2=0"

"D=1600-4*2=1592"

"Q_1=(40-\sqrt{1592})/2=-0.05" (it cannot be negative)

"Q_2=(40+\sqrt{1592})/2=39.94"

So, the profit maximizing level of production is 39.94